What is the annual yield for an investment that earns 10% compounded semi-annually? Please include the formula that you used to solve this problem.
Answer by stanbon(75887)

What is the annual yield for an investment that earns 10% compounded semi-annually? Please include the formula that you used to solve this problem.
----------------
Formula:: A(t) = P(1+(r/n))^(n*t)
--------------
A(1) = P(1+(0.10/2)]^(2*1) = P*1.05^2
-------
Cheers,
Stan H.
---------